# Spatial Context-based Self-Supervised Learning for Handwritten Text Recognition

### v1

@@ Appendix A Data augmentation
iaa.Sequential([
iaa.SomeOf((1, 5), [
iaa.LinearContrast((0.5, 1.0)),
iaa.GaussianBlur((0.5, 1.5)),
iaa.Sharpen(alpha=(0.0, 0.5), lightness=(0.0, 0.5)),
iaa.PiecewiseAffine(scale=(0.02, 0.03), mode=’edge’),
iaa.PerspectiveTransform(scale=(0.01, 0.02)),
], random_order=True)]
@@
Furthermore, drawing on established practices in the literature, we also employ a soft data augmentation strategy during the supervised stage. This includes subtle cropping, linear contrast adjustments, and Gaussian blur.
Appendix B Detailed results
Owing to the extensive experimentation carried out in this paper, this section offers a complete compilation of results for each method with respect to the specific architectures employed and the datasets used for pre-training and evaluation.
In Tab. B we report the quality results in terms of the CER and WAcc metrics.
In Tab. B we provide the semi-supervised results in terms of WAcc and Single Edit Distance (ED1) metrics. ED1 expresses word-level accuracy up to a single edit operation. Note that, the exclusion of the latter metric in the manuscript was primarily due to concerns regarding clarity. Nevertheless, in this section, we opted to include it to enhance the comprehensiveness of the results presented. It is worth highlighting that the incorporation of this metric does not deviate from the conclusions drawn in the paper; instead, it serves to reinforce and bolster them.
CER WAcc
IAM CVL G.W. IAM CVL G.W.
Geometric
Rotation 27.4 19.7 35.1 53.6 56.8 25.5
Flip 23.8 20.5 28.8 56.1 57.3 43.8
[…]
Jigsaw 52.4 28.2 31.3 31.6 46.0 42.2
Sorting 79.7 53.0 31.4 19.4 12.3 43.2
ChaCo 25.2 23.7 84.7 48.3 47.0 04.7
(a) Representation quality under CTC decoder.
CER WAcc
IAM CVL G.W. IAM CVL G.W.
Geometric
Rotation 19.0 11.9 22.6 64.6 79.2 59.8
Flip 17.4 15.3 19.3 64.7 75.1 61.1
[…]
Jigsaw 49.1 19.6 38.3 35.7 70.5 44.9
Sorting 72.1 58.8 64.0 8.6 24.7 19.8
ChaCo 12.2 11.9 69.2 71.6 80.2 10.0
(b) Representation quality under TD decoder.
(c) Representation quality. The frozen part is only the visual encoder pre-trained in the SSL method. Bold values indicate the best results under the same conditions of pre-training and test.
WAcc ED1
IAM CVL G.W. IAM CVL G.W.
5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100%
Baseline
No pretrain 26.5 42.2 70.4 07.5 14.2 69.7 13.5 16.0 78.5 42.1 64.9 88.1 17.4 25.4 85.6 26.3 44.6 92.3
ChaCo 30.0 46.3 64.6 13.1 13.2 67.4 15.7 21.2 72.6 47.7 70.2 83.9 34.6 40.7 84.0 28.7 34.9 89.6
[…]
Rotation 42.6 48.9 67.7 10.0 16.4 70.1 17.0 36.0 82.4 67.7 74.4 88.2 27.2 38.8 85.9 32.9 56.9 94.0
Flip 42.1 50.6 72.1 17.1 35.0 71.1 18.8 32.5 79.1 68.4 75.6 89.4 38.7 61.1 86.8 35.9 54.3 93.2
Jigsaw 42.4 51.4 67.7 05.3 12.8 55.4 12.9 19.5 72.9 66.8 75.2 86.7 18.1 31.2 77.3 26.4 43.9 89.6
Sorting 23.9 41.6 61.1 09.3 14.7 70.0 13.8 08.9 76.6 38.1 63.9 81.1 19.3 26.2 84.7 24.1 33.1 91.4
(d) Semi-supervised results under CTC decoder.
WAcc ED1
IAM CVL G.W. IAM CVL G.W.
5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100% 5% 10% 100%
Baseline
No pretrain 10.3 50.3 74.8 05.9 14.2 68.7 06.3 02.1 71.3 21.7 63.6 87.6 07.9 25.4 85.6 09.2 02.4 82.8
ChaCo 47.3 53.8 75.1 21.2 38.0 83.4 05.9 07.0 73.2 66.6 74.0 88.2 27.1 44.9 87.9 06.6 08.2 83.1
[…]
Rotation 47.0 54.5 74.5 21.5 51.8 85.5 03.6 06.9 78.5 65.6 74.8 87.2 28.3 61.4 90.6 07.7 08.7 89.5
Flip 43.8 54.3 74.6 25.2 53.1 86.1 06.0 11.0 78.0 60.2 72.0 88.1 29.7 62.8 90.7 08.9 13.5 90.3
Jigsaw 43.5 51.7 74.3 21.5 38.1 84.8 14.6 20.2 77.1 60.5 69.1 87.1 27.8 44.6 88.5 18.9 26.9 88.7
Sorting 25.2 39.8 75.2 05.4 22.0 83.6 11.9 14.2 76.1 34.5 49.4 87.2 07.0 26.5 87.0 15.4 17.2 86.6
(e) Semi-supervised results under TD decoder.
(f) Semi-supervised results. The models pre-trained with the SSL methods are fine-tuned with 5%, 10% and 100% of the labelled data. Bold values indicate the best results under the same conditions of pre-training and test.

### v2

@@ A.1 Data augmentation
Furthermore, drawing on established practices in the literature, we also employ the data augmentation strategy during the supervised stage. This includes the previous data augmentations excluding hard geometric augmentations.
A.2 Visual encoder architectural details
In this subsection, Table 5 details the precise configuration of the encoder architecture contemplated in the work. Note that, as commented in the manuscript, this scheme stands for a ResNet-29 architecture [9] with slight modifications.
Table 5: Architectural details of the neural encoder considered. Block refers to a block of convolutional layers.
Layer Configuraton
Kernel, Channels Stride Padding
Convolution $3\times 3$, 32 (1, 1) same
Convolution $3\times 3$, 64 (1, 1) same
Max pooling $2\times 2$, - (2, 2) (0, 0)
Block $\begin{bmatrix}3\times 3,&128\\ 3\times 3,&128\end{bmatrix}$$\times 1$ (1, 1) same
Max pooling $2\times 2$, - (2, 2) (0, 0)
Block $\begin{bmatrix}3\times 3,&256\\ 3\times 3,&256\end{bmatrix}$$\times 2$ (1, 1) same
Max pooling $2\times 2$, - (2, 2) (0, 0)
Block $\begin{bmatrix}3\times 3,&512\\ 3\times 3,&512\end{bmatrix}$$\times 5$ (1, 1) same
Convolution $3\times 3$, 512 (1, 1) same
Max pooling $2\times 2$, - (2, 2) (0, 0)
Block $\begin{bmatrix}3\times 3,&512\\ 3\times 3,&512\end{bmatrix}$$\times 3$ (1, 1) same
Max pooling $2\times 2$, - (2, 2) (0, 0)
Convolution $3\times 3$, 512 (1, 1) same
A.3 RIMES partition details
In this work, we utilized the latest version of the RIMES word dataset [15]. This dataset is organized into 57 folders, each referred to as a block, containing each one 100 letters represented as word-level images. To our best knowledge, there is no official partitioning scheme for this version of the dataset, nor are there prior studies that use it. Consequently, we have created arbitrary partitions for this work whose details are now provided to ensure reproducibility and facilitate future comparisons:
Table 6: Table showing the partitions employed in RIMES dataset
Partition Blocks Samples
Train 1-34 139490
Validation 35-39 20217
Test 40-57 68651